Prepare an amortization table for a house you would like to purchase. Prepare it for monthly for two years of the 30 years. Decide how much you can afford knowing it should be 1/3 of your income. Your income is 268,000. You have 30 years and 5% interest. How would that change if you choose biweekly payments instead of monthly. Give an example of each. Create another spreadsheet in the same workbook to show your answer.
Answer #1

Due to character limit I am attaching the image for amortization schedule of biweekly payment.

Amortization Schedule A B C=A*5%/12 D=B-C E=A-D Income               268,000.00
Monthly Payments Beginning Amount Paid Interest Owed Principal paid Ending I can afford 1/3rd
1 16,641,157.17        89,333.33         69,338.15      19,995.18 16,621,161.99 i.e. monthly payment will be 89,333.33 268000/3
2 16,621,161.99        89,333.33         69,254.84      20,078.49 16,601,083.50
3 16,601,083.50        89,333.33         69,171.18      20,162.15 16,580,921.35 PMT= PV*i
4 16,580,921.35        89,333.33         69,087.17      20,246.16 16,560,675.19 1-(1+i)^-n
5 16,560,675.19        89,333.33         69,002.81      20,330.52 16,540,344.67
6 16,540,344.67        89,333.33         68,918.10      20,415.23 16,519,929.43 PMT= Monthly payment 89,333.33
7 16,519,929.43        89,333.33         68,833.04      20,500.29 16,499,429.14 PV= Loan amount ?
8 16,499,429.14        89,333.33         68,747.62      20,585.71 16,478,843.43 n= no. of payments 30*12=        360.00
9 16,478,843.43        89,333.33         68,661.85      20,671.49 16,458,171.94 i 5%/12 0.0041667
10 16,458,171.94        89,333.33         68,575.72      20,757.62 16,437,414.33
11 16,437,414.33        89,333.33         68,489.23      20,844.11 16,416,570.22
12 16,416,570.22        89,333.33         68,402.38      20,930.96 16,395,639.26 89333.33 PV*0.00416666666666667
13 16,395,639.26        89,333.33         68,315.16      21,018.17 16,374,621.09 1-(1+0.00416666666666667)^-360
14 16,374,621.09        89,333.33         68,227.59      21,105.75 16,353,515.35
15 16,353,515.35        89,333.33         68,139.65      21,193.69 16,332,321.66 89333.33 PV
16 16,332,321.66        89,333.33         68,051.34      21,281.99 16,311,039.67 0.004166667 1-(1.00416666666666667)^-360
17 16,311,039.67        89,333.33         67,962.67      21,370.67 16,289,669.00
18 16,289,669.00        89,333.33         67,873.62      21,459.71 16,268,209.29 21439999.2 PV
19 16,268,209.29        89,333.33         67,784.21      21,549.13 16,246,660.16 1-0.223826595641891
20 16,246,660.16        89,333.33         67,694.42      21,638.92 16,225,021.24
21 16,225,021.24        89,333.33         67,604.26      21,729.08 16,203,292.16 21439999.2 PV
22 16,203,292.16        89,333.33         67,513.72      21,819.62 16,181,472.55 0.776173404 0.7761734
23 16,181,472.55        89,333.33         67,422.80      21,910.53 16,159,562.02
24 16,159,562.02        89,333.33         67,331.51      22,001.82 16,137,560.19 PV 21439999.2*0.776173404358109
1,640,403.02 PV         16,641,157.17 Loan Value
Point to Note:
Loan value will be greater by $ 1,395,468.97 if bi weekly payment plan adopted.
Interest paid for first 2 year will be more by $ 136,973.11 if bi weekly payment plan adopted.